# Speak, Segment, Track, Navigate: An Interactive System for Video-Guided Skull-Base Surgery

### 17 Mar 2026 v1

@@ III-F Spatial Pose Tracking
Restricting inference to the surgical foreground concentrates depth estimation on geometrically meaningful regions and reduces variability introduced by non-surgical scene elements. By limiting the input to tool and anatomical structures, the depth network focuses on regions directly involved in downstream geometric reasoning, resulting in improved depth consistency and more stable scale alignment.
To resolve the scale ambiguity of monocular depth, we estimate an affine mapping
@@
Figure 7: Foreground-conditioned monocular depth scaling for metric tool reconstruction. From the real-time video stream, the mask tracking module produces a tool mask and a foreground mask, which are used to suppress background appearance and form a foreground-masked frame. The registered anatomy provides a reliable metric depth reference, enabling an affine scale alignment of DepthAnything v2 predictions to obtain a scaled depth map in millimeters for both anatomy and tool regions.
The parameters $(\alpha,\beta)$ are then applied to the full depth map, yielding dense metric depth for both anatomy and tool regions. Using the tracked tool mask $M_{f}^{\mathrm{tool}}(\mathbf{u})\in\{0,1\}$, we back-project the tool region into a camera-frame point cloud. Let $Z_{f}(\mathbf{u})$ denote the scaled metric depth at pixel $\mathbf{u}=(u,v)^{\top}$. Using the camera intrinsics, we lift each tool pixel to a 3D camera-frame point via
$\mathbf{X}^{C}(\mathbf{u})=\begin{bmatrix}(u-c_{x})Z_{f}(\mathbf{u})/f_{x}\\ (v-c_{y})Z_{f}(\mathbf{u})/f_{y}\\ Z_{f}(\mathbf{u})\end{bmatrix}$.
The tool point set is then
$\mathcal{P}^{C}_{\mathrm{tool}}=\left\{\mathbf{X}^{C}(\mathbf{u})\;\middle|\;M_{f}^{\mathrm{tool}}(\mathbf{u})=1\right\}.$
We estimate a coarse 3D tool-axis prior $\tilde{\mathbf{d}}^{C}_{3\mathrm{D}}\in\mathbb{S}^{2}$, where $\mathbb{S}^{2}=\{\mathbf{x}\in\mathbb{R}^{3}:\|\mathbf{x}\|_{2}=1\}$ denotes the unit sphere, by principal component analysis (PCA) on $\mathcal{P}^{C}_{\mathrm{tool}}$, using the first principal component as the dominant direction. Because this estimate is affected by depth noise we treat $\tilde{\mathbf{d}}^{C}_{3\mathrm{D}}$ only as an initial geometric prior and refine it using mesh- and mask-derived constraints in the pose initialization step below.
Pose Initialization
[…]
Camera model: We use the pinhole projection $\pi(\cdot;K):\mathbb{R}^{3}\!\to\!\mathbb{R}^{2}$ with intrinsics $K=\{f_{x},f_{y},c_{x},c_{y}\}$:
$\pi\!\left(\begin{bmatrix}X\\ Y\\ Z\end{bmatrix};K\right)=\begin{bmatrix}f_{x}X/Z+c_{x}\\ f_{y}Y/Z+c_{y}\end{bmatrix}$.
[…]
@@
    Length-induced tilt constraint: $r=\ell_{\mathrm{mask}}^{\mathrm{px}}/\ell_{\mathrm{mesh}}^{\mathrm{px}}$, where $\ell_{\mathrm{mesh}}^{\mathrm{px}}$ is the projected CAD length under the provisional axis, constraining out-of-plane tilt.
[…]
Given the camera-frame tip $\mathbf{p}^{C}_{\mathrm{tip}}$ and the provisional axis $\tilde{\mathbf{d}}^{C}_{3\mathrm{D}}$, we form the length-$L$ segment
$\mathbf{P}_{0}=\mathbf{p}^{C}_{\mathrm{tip}},\qquad\mathbf{P}_{1}=\mathbf{p}^{C}_{\mathrm{tip}}+L\,\tilde{\mathbf{d}}^{C}_{3\mathrm{D}},$
and define the expected pixel length
$\ell_{\mathrm{mesh}}^{\mathrm{px}}=\left\|\pi(\mathbf{P}_{1};K)-\pi(\mathbf{P}_{0};K)\right\|_{2}.$
Let $\ell_{\mathrm{mask}}^{\mathrm{px}}$ denote the observed 2D tool length from mask PCA, and define
$r=\frac{\ell_{\mathrm{mask}}^{\mathrm{px}}}{\ell_{\mathrm{mesh}}^{\mathrm{px}}},\qquad r\leftarrow\mathrm{clip}\!\left(r,\,r_{\min},\,r_{\max}\right).$
2D alignment constraint at the tip: Let $\mathbf{p}^{C}_{\mathrm{tip}}=(X,Y,Z)^{\top}$ and define normalized image coordinates $x=X/Z$ and $y=Y/Z$. For a candidate direction $\mathbf{d}^{C}=(d_{x},d_{y},d_{z})^{\top}$, consider the 3D line through the tip
@@
We enforce image-plane orientation consistency by requiring
$\mathbf{g}(\mathbf{d}^{C})\parallel\mathbf{d}^{C}_{2\mathrm{D}},$
so that the projected 3D axis matches the mask PCA direction.
Length-induced tilt constraint. For a unit direction $\mathbf{d}^{C}=(d_{x},d_{y},d_{z})^{\top}$, the image-plane magnitude
@@
The out-of-plane component follows from the unit-norm constraint $\|\mathbf{d}^{C}\|_{2}=1$:
$d_{z}=\pm\sqrt{1-\rho^{2}}.$
The two signs correspond to the tool axis pointing into the camera (toward the imaging sensor) or out of the camera (away from the sensor). We disambiguate this by selecting the sign that is consistent with the DA2-derived 3D prior $\tilde{\mathbf{d}}^{C}_{3\mathrm{D}}$:
$d_{z}=\mathrm{sign}(\tilde{d}_{z})\sqrt{1-\rho^{2}}.$
Closed-form solution: Let $\mathbf{d}^{C}_{2\mathrm{D}}=(d_{2x},d_{2y})^{\top}$ be the unit mask PCA direction. The constraint $\mathbf{g}(\mathbf{d}^{C})\parallel\mathbf{d}^{C}_{2\mathrm{D}}$ implies that $\mathbf{g}(\mathbf{d}^{C})$ must lie on the 1D subspace spanned by $\mathbf{d}^{C}_{2\mathrm{D}}$, i.e., there exists a scalar $\alpha\in\mathbb{R}$ (the signed image-plane scale along $\mathbf{d}^{C}_{2\mathrm{D}}$) such that
$\mathbf{g}(\mathbf{d}^{C})=\alpha\,\mathbf{d}^{C}_{2\mathrm{D}}.$
and $\mathbf{d}^{C}_{2\mathrm{D}}=(d^{C}_{2\mathrm{x}},d^{C}_{2\mathrm{y}})^{\top}$ Substituting the definition of $\mathbf{g}(\mathbf{d}^{C})$ gives
$f_{x}(d_{x}-xd_{z})=\alpha d_{2x},\qquad f_{y}(d_{y}-yd_{z})=\alpha d_{2y},$
@@
$d_{x}=xd_{z}+\alpha\frac{d_{2x}}{f_{x}},\qquad d_{y}=yd_{z}+\alpha\frac{d_{2y}}{f_{y}}.$
@@
$\left(xd_{z}+\alpha\frac{d_{2x}}{f_{x}}\right)^{2}+\left(yd_{z}+\alpha\frac{d_{2y}}{f_{y}}\right)^{2}=\rho^{2}.$
@@
$A=\frac{d_{2x}}{f_{x}},\qquad B=\frac{d_{2y}}{f_{y}},\qquad C_{x}=xd_{z},\qquad C_{y}=yd_{z}.$
@@
$\mathbf{d}^{C}=\mathbf{d}^{C}_{(k^{*})}.$
Since $a=A^{2}+B^{2}$ with $A=d_{2x}/f_{x}$ and $B=d_{2y}/f_{y}$ ($f_{x},f_{y}>0$), we have $a>0$ whenever the 2D PCA direction $\mathbf{d}^{C}_{2\mathrm{D}}$ is valid (nonzero). Although PCA is sign-ambiguous, sign flips do not affect $a$ because it depends only on squared terms. Thus $a\approx 0$ indicates a degenerate 2D estimate; if $a\approx 0$ or $\Delta<0$, we fall back to the DA2-derived 3D prior $\tilde{\mathbf{d}}^{C}_{3\mathrm{D}}$.
Beyond numerical validity, stability depends on the reliability of the length ratio
$r=\frac{\ell_{\mathrm{mask}}^{\mathrm{px}}}{\ell_{\mathrm{mesh}}^{\mathrm{px}}}.$
We therefore instruct the user to briefly hold the tool such that the entire shaft is visible before initialization; in this configuration, $\ell_{\mathrm{mask}}^{\mathrm{px}}$ more faithfully reflects the physical length, making the tilt magnitude inferred from $r$ stable and consistent with the mesh projection.
[…]
Rodrigues’ formula gives
The resulting rigid transform $T^{C}_{\mathrm{mesh}}=(R_{C},\mathbf{t}_{C})$ initializes the tool pose in the camera frame. We then render the posed mesh to obtain a tool-only depth image $Z_{\mathrm{tool}}$ (mm) and its silhouette. Using the same pixel-to-3D lifting as in the metric reconstruction step, the rendered depth is converted to a camera-frame point set
$\mathbf{X}^{C}(\mathbf{u};Z_{\mathrm{tool}})=\begin{bmatrix}(u-c_{x})Z_{\mathrm{tool}}(\mathbf{u})/f_{x}\\ (v-c_{y})Z_{\mathrm{tool}}(\mathbf{u})/f_{y}\\ Z_{\mathrm{tool}}(\mathbf{u})\end{bmatrix}$
$\mathcal{P}_{\mathrm{tool}}^{C}=\left\{\mathbf{X}^{C}(\mathbf{u};Z_{\mathrm{tool}})\;\middle|\;\mathbf{u}\in\Omega_{\mathrm{tool}}\right\}.$
@@
After initialization, the observed 2D mask length $\ell_{\mathrm{mask}}^{\mathrm{px}}(t)$ is not guaranteed to reflect the true projected tool extent. Changes in $\ell_{\mathrm{mask}}^{\mathrm{px}}$ may result from genuine out-of-plane orientation changes (either shortening or lengthening due to perspective), but may also arise from partial occlusion or truncation at the image boundary. Because the tracker primarily provides 2D skeleton cues (tip location and principal image-axis direction), the 2D signal alone cannot reliably distinguish geometric tilt variation from visibility artifacts.
At each frame $t$, we form two temporally consistent pose proposals derived from the previous-frame estimate:
@@
Figure 9: Comparison of pose hypotheses used for tracking. From left to right: 2D tool mask ground truth, the tilt-adjusted geometry propagation hypothesis, and the no-tilt geometry propagation hypothesis. The hypothesis with higher mask overlap (F1 score) is selected to ensure robust tracking under partial occlusion and truncation
@@
Figure 10: Comparison of silhouette hypotheses used for pose tracking. From left to right: 2D tool mask ground truth, the Depth Anything v2 (DA2)-based silhouette measurement, and the temporally propagated geometry silhouette. The hypothesis with higher mask overlap (F1 score) is selected to ensure robust tracking under occlusion and rapid viewpoint changes.
Because both hypotheses are derived from the previous pose, tracking relies strongly on temporal consistency. To improve robustness, we additionally retain an independent fallback direction obtained from per-frame 3D PCA on the scaled Depth Anything v2 (DA2) point cloud. Let $\hat{M}_{\mathrm{DA2}}(t)$ denote the silhouette obtained from this direction. The final pose at frame $t$ is selected by maximizing
$\mathrm{F1}\big(\hat{M}_{k}(t),\hat{M}_{\mathrm{DA2}}(t)\big)$
$k\in\{\text{tilt, no-tilt}\}.$
In this formulation, temporal consistency governs nominal geometry propagation tracking, while the DA2-based measurements provides a geometrically grounded recovery mechanism when temporal assumptions fail.
@@ IV Experiments and Results
We evaluated the proposed framework in video-guided skull base surgery scenarios using an ex vivo skull-base setup. The surgical drill was tracked concurrently using (i) the optical tracker and (ii) the proposed vision-based perception and pose-tracking pipeline. To enable direct comparison, we time-matched the two data streams and expressed both pose estimates in the same anatomy coordinate frame $\mathcal{F}_{A}$ obtained from the registration procedure.
Across three trials, the user performed continuous tool motion including approach, drilling-like contact motion, and withdrawal. All reported statistics are computed over time-matched frames after initialization and registration, excluding frames where either system produced invalid measurements (e.g., missing optical tracking due to marker dropout).
IV-A Pose Tracking Accuracy
Tool-tip translation error
For each time-matched frame $i$, we obtained the vision tip position $\mathbf{p}^{V}_{A,i}\in\mathbb{R}^{3}$ by transforming the estimated camera-frame tip into $\mathcal{F}_{A}$, and we obtained the optical-tracker tip position $\mathbf{p}^{O}_{A,i}\in\mathbb{R}^{3}$ by applying the measured marker pose and a fixed pivot offset from the marker body to the drill tip. The signed position error in the anatomy frame was computed as
$\Delta\mathbf{p}_{i}=\mathbf{p}^{V}_{A,i}-\mathbf{p}^{O}_{A,i}.$ (3)

### 16 Apr 2026 v2

@@ III-G Spatial Pose Tracking
To resolve the scale ambiguity of monocular depth estimation, we estimate an affine mapping
@@
Figure 7: Foreground-conditioned monocular depth scaling for metric tool reconstruction. From the real-time video stream, the mask tracking module enables a foreground-masked frame. The registered anatomy provides a reliable metric depth reference, enabling an affine scale alignment of DepthAnything v2 predictions to obtain a scaled depth map in millimeters for both anatomy and tool regions.
The parameters $(\alpha,\beta)$ are then applied to the full depth map, yielding dense metric depth for both anatomy and tool regions. Using the tracked tool mask $M_{f}^{\mathrm{tool}}(\mathbf{u})\in\{0,1\}$, we back-project the tool region into a camera-frame $C$ point cloud $\mathcal{P}^{C}_{\mathrm{tool}}$ using intrinsics. The tool’s principle direction is extracted via:
$\mathbf{d}^{C}=\text{PCA($\mathcal{P}^{C}_{\mathrm{tool}}$)}$
Rigid alignment of the CAD mesh
Once the camera-frame axis $\mathbf{d}^{C}$ has been determined, we initialize the mesh pose by aligning its longitudinal axis $\mathbf{a}_{\mathrm{local}}$ to $\mathbf{d}^{C}$ and anchoring the mesh tip $\mathbf{p}^{\mathrm{mesh}}_{\mathrm{tip}}$ at the depth-derived tip location $\mathbf{p}^{C}_{tip}$.We compute compute a rotation $R_{C}\in SO(3)$ that maps $\mathbf{a}_{\mathrm{local}}$ to $\mathbf{d}^{C}$ with Rodrigues’ formula such that:
The resulting rigid transform $T^{C}_{\mathrm{mesh}}=(R_{C},\mathbf{t}_{C})$ transforms the tool from its mesh local coordinate to the camera frame pose that matches the latest position and orientation of the surgical tool.
However, this approach suffers from noise in the relative depth estimation, which results in fluctuations and inaccuracies in the pose estimation. For this reason, we designed the hybrid approach which uses the $\mathbf{d}^{C}$ derived from monocular depth estimation as an initial geometric prior and refine it using mask-derived constraints in the steps below.
Hybrid Pose Tracking Approach
Figure 8: 3D tool pose initialization from mask geometry and metric depth. From the tracked tool mask and foreground-conditioned metric depth, the 2D branch extracts the principal direction $\mathbf{d}_{M}^{C}$ and apparent length $\ell^{px}_{mask}$, while the 3D branch back-projects the masked depth to obtain the tool point set $\mathcal{P}^{C}_{\mathrm{tool}}$, a coarse PCA axis prior $\tilde{\mathbf{d}}^{C}_{3D}$, and the tip position $\mathbf{p}^{C}_{\mathrm{tip}}$. From the CAD model, we use the mesh tip $\mathbf{p}^{\mathrm{mesh}}_{\mathrm{tip}}$, tool length $L$, and canonical local axis $\mathbf{a}_{\mathrm{local}}$. The pose solver computes the ratio $r=\ell^{px}_{mask}/\ell^{px}_{mesh}$, refines the camera-frame axis $\mathbf{d}^{C}_{3D}$ using the 2D projection constraint and coarse 3D prior, and aligns the mesh by $(R(\mathbf{d}^{C}_{3D}),\mathbf{t}^{C})$ to recover $T^{C}_{\mathrm{mesh}}$, which is then rendered as the tool overlay.
This approach treats pose initialization and tracking as two components with separate 2D constraints to enforce temporal consistency. Camera model: We use the pinhole projection $\pi(\cdot;K):\mathbb{R}^{3}\!\to\!\mathbb{R}^{2}$ with intrinsics $K=\{f_{x},f_{y},c_{x},c_{y}\}$. Notation: In the previous formulation, $\mathbf{d}^{C}$ denoted the camera-frame direction. In the present formulation, however, this quantity is redefined as a coarse depth-based prior obtained from Depth Anything. To avoid ambiguity, we denote this prior by $\tilde{\mathbf{d}}^{C}_{3\mathrm{D}}$. Tip Depth Extraction: Instead of using tip depth derived from scaled monocular depth estimation, we directly use depth from the registered anatomy to improve temporal consistencies.
Pose Initialization: We will estimate a refined unit direction $\mathbf{d}^{C}$ using:
@@
    Length-induced pitch constraint: $r=\ell_{\mathrm{mask}}^{\mathrm{px}}/\ell_{\mathrm{mesh}}^{\mathrm{px}}$, where $\ell_{\mathrm{mesh}}^{\mathrm{px}}$ is the projected CAD length under the provisional axis, constraining out-of-plane tilt.
Enforcing Yaw Constraint with 3D Prior: For a refined direction $\mathbf{d}^{C}$, consider the 3D line through the tip:
@@
Let $\mathbf{d}^{C}_{Mask}=(d_{x}^{M},d_{y}^{M})^{\top}$ be the unit mask PCA direction. We enforce image-plane orientation consistency by requiring
$\mathbf{g}(\mathbf{d}^{C})\parallel\mathbf{d}^{C}_{Mask},$
Length-induced pitch constraint: For a unit direction $\mathbf{d}^{C}=(d_{x},d_{y},d_{z})^{\top}$, the image-plane magnitude
@@
Closed-form solution: The constraint $\mathbf{g}(\mathbf{d}^{C})\parallel\mathbf{d}^{C}_{\text{Mask}}$ implies that $\mathbf{g}(\mathbf{d}^{C})$ must lie on the 1D subspace spanned by $\mathbf{d}^{C}_{2\mathrm{D}}$, i.e., there exists a scalar $\alpha\in\mathbb{R}$ (the signed image-plane scale along $\mathbf{d}^{C}_{2\mathrm{D}}$) such that
$\mathbf{g}(\mathbf{d}^{C})=\alpha\,\mathbf{d}^{C}_{\text{Mask}}.$
Substituting the definition of $\mathbf{g}(\mathbf{d}^{C})$ gives
$f_{x}(d_{x}-xd_{z})=\alpha d_{x}^{M},\qquad f_{y}(d_{y}-yd_{z})=\alpha d_{y}^{M},$
@@
$d_{x}=xd_{z}+\alpha\frac{d_{x}^{M}}{f_{x}},\qquad d_{y}=yd_{z}+\alpha\frac{d_{y}^{M}}{f_{y}}.$
@@
$\left(xd_{z}+\alpha\frac{d_{x}^{M}}{f_{x}}\right)^{2}+\left(yd_{z}+\alpha\frac{d_{y}^{M}}{f_{y}}\right)^{2}=\rho^{2}.$
@@
$A=\frac{d_{x}^{M}}{f_{x}},\qquad B=\frac{d_{y}^{M}}{f_{y}},\qquad C_{x}=xd_{z},\qquad C_{y}=yd_{z}.$
@@
$\boxed{\mathbf{d}^{C}=\mathbf{d}^{C}_{(k^{*})}}$
Since $a=A^{2}+B^{2}$ with $A=d_{2x}/f_{x}$ and $B=d_{2y}/f_{y}$ ($f_{x},f_{y}>0$), we have $a>0$ whenever the 2D PCA direction $\mathbf{d}^{C}_{2\mathrm{D}}$ is valid (nonzero). Thus, this method would only fail when the tool disappears from the scene or is in extreme angles untraceable by the mask tracking model.
With the previously discussed rigid alignment process, we derive the rigid transform $T^{C}_{\mathrm{mesh}}=(R_{C},\mathbf{t}_{C})$ which initializes the tool pose in the camera frame. We then render the posed mesh to obtain a tool-only depth image $Z_{\mathrm{mesh}}$ and its silhouette. Using the same pixel-to-3D computation as in the metric reconstruction step, the rendered depth is converted to a camera-frame point set $\mathcal{P}_{\mathrm{mesh}}^{C}$.
@@
After the initialization, 2D constraints will remain unchanged, but the pitch of the tool will be updated using the observed cross-frame 2D tool mask length changes. However, the limitation with this approach is that the changes in $\ell_{\mathrm{mask}}^{\mathrm{px}}$ may result from genuine out-of-plane orientation changes (either shortening or lengthening due to perspective), but may also arise from partial occlusion or truncation at the image boundary. Because the tracker primarily provides 2D skeleton cues (tip location and principal image-axis direction), the 2D signal alone cannot reliably distinguish geometric tilt variation from visibility artifacts.
Therefore, at each frame $t$, we form two temporally consistent pose proposals derived from the previous-frame estimate:
@@
Algorithm 1 Hybrid Tool Pose Tracking
1:Inputs:
2:   $I_{t}$: RGB image at frame $t$
3:   $M_{t}^{\mathrm{tool}}$: binary mask of the target tool
4:   $M_{t}^{\mathrm{anat}}$: binary mask(s) of registered anatomy
5:   $K$: camera intrinsic matrix
6:   $S_{t}^{\mathrm{reg}}$: registration-based metric depth prior
7:   $T^{C}_{\mathrm{mesh}}(t\!-\!1)$: pose in the camera frame from frame t1t\!-\!1
8:Output:
9:   $T^{C}_{\mathrm{mesh}}(t)$: updated tool mesh pose at $t$
10:
[…]
18:
19:if $t=1$ then $\triangleright$ initialize pose from depth and mask geometry
20:  return $\mathrm{InitPose}\!\left(\tilde{\mathbf{d}}^{C}_{3\mathrm{D}},\mathbf{d}^{C}_{2\mathrm{D}},\mathbf{p}^{C}_{\mathrm{tip}}\right)$
21:end if
22:
[…]
27:
28:return $T^{C}_{\mathrm{mesh}}(t)$
@@ IV Experiments and Results
We evaluated the proposed framework in video-guided skull base surgery scenarios using an ex vivo skull-base setup. The experiments were designed to assess three aspects of the system: (1) the accuracy of anatomy registration, (2) the accuracy and temporal stability of tool pose tracking relative to an optical tracking reference, and (3) the efficiency of the overall interactive workflow.
During each trial, the surgical drill was tracked (wrt. anatomical coordinate frame) concurrently using (i) the optical tracker and (ii) the proposed vision-based perception and pose-tracking pipeline. To enable direct comparison, the two data streams were time-synchronized and both pose estimates were expressed in the same camera coordinate frame $\mathcal{F}_{C}$ through transformations obtained from the registration procedure.
Across three trials, the user performed continuous tool motion including approach, drilling-like contact motion, and withdrawal. All reported statistics are computed over time-synchronized frames after initialization and registration, excluding frames when optical tracking system did not provide a valid measurement due to marker occlusion or temporary loss of line-of-sight.
IV-A Registration Accuracy
We evaluated the anatomy-to-camera registration pipeline using reprojection root-mean-square error (RMSE). Two interaction modes were compared: manual clicking and the proposed virtual cursor interface. Manual clicking involves direct point selection, while the virtual cursor involves landmark selection using the tracked tool tip. In both cases, 2D image points were associated with known 3D anatomical landmarks and used to estimate camera pose via PnP.
Overall, the virtual cursor achieved sub-mm registration accuracy and remained close to the manual-clicking baseline in two of the three trials. The higher error observed in Trial 3 reflects the sensitivity of virtual cursor interaction to accumulated tip-localization error during landmark selection. Nevertheless, the mean error remained within a practically useful range for downstream guidance tasks.
TABLE I: Registration pipeline RMSE for manual clicking and virtual cursor interaction.
Manual Clicking Virtual Cursor
Trial 1 3.43 px (0.22mm)(0.22\ $\mathrm${mm}) 3.63 px (0.23mm)(0.23\ $\mathrm${mm})
Trial 2 3.03 px (0.18mm)(0.18\ $\mathrm${mm}) 2.48 px (0.15mm)(0.15\ $\mathrm${mm})
Trial 3 2.28 px (0.15mm)(0.15\ $\mathrm${mm}) 11.64 px (0.74mm)(0.74\ $\mathrm${mm})
Mean 2.91 px (0.18mm)(0.18\ $\mathrm${mm}) 5.92 px (0.37mm)(0.37\ $\mathrm${mm})
IV-B Pose Tracking Accuracy
We evaluated the vision-based pose tracking module against the optical tracking system. We report both translation error and rotation errors.
Tool-tip translation error. For each time-synchronized frame $i$, we obtained the vision-based tip position $\mathbf{p}^{V}_{C,i}\in\mathbb{R}^{3}$ and the optical-tracker tip position $\mathbf{p}^{O}_{C,i}\in\mathbb{R}^{3}$, both expressed in the camera frame. The signed position error was computed as
$\Delta\mathbf{p}_{i}=\mathbf{p}^{V}_{C,i}-\mathbf{p}^{O}_{C,i}.$ (3)